OpenStudy (anonymous):

Solve the initial value problem y''+9y=0 when y(0)=2 and y'(0)=3. Can't seem to get the correct answer, help would be appreciated

OpenStudy (amistre64):

what is your characteristic equation?

OpenStudy (anonymous):

well it wil be y=c1*e^(m1*x)+c2e^(m2*x). The roots are m1=0 and m2=-9, but I can't get the correct values for c1 and c2...

OpenStudy (amistre64):

your roots are wrong

OpenStudy (amistre64):

m^2 + 9 = 0 m^2 = -9 m = 0 +- 3i

OpenStudy (anonymous):

ah, missed a sign somewhere

OpenStudy (amistre64):

y = e^0 (c1 cos(3x) + c2 sin(3x)) y = c1 cos(3x) + c2 sin(3x)

OpenStudy (anonymous):

Got the rest of it , thanks a bunch